# ScarFinder: a detector of optimal scar trajectories in quantum many-body dynamics

### v1

@@ Abstract
Mechanisms that give rise to coherent quantum dynamics, such as quantum many-body scars, have recently attracted much interest as a way of controlling quantum chaos. However, identifying the presence of quantum scars in general many-body Hamiltonians remains an outstanding challenge. Here we introduce ScarFinder, a variational framework that reveals possible scar-like dynamics without prior knowledge of scar states or their algebraic structure. By iteratively evolving and projecting states within a low-entanglement variational manifold, ScarFinder isolates scarred trajectories by suppressing thermal contributions. We validate the method on the analytically tractable spin-1 XY model, recovering the known scar dynamics, as well as the mixed field Ising model, where we capture and generalize the initial conditions previously associated with “weak thermalization”. We then apply the method to the PXP model of Rydberg atom arrays, efficiently characterizing its mixed phase space and finding a previously unknown trajectory with nearly-perfect revival dynamics in the thermodynamic limit. Our results establish ScarFinder as a powerful, model-agnostic tool for identifying and optimizing coherent dynamics in quantum many-body systems.
@@ I Introduction
Recent progress in quantum simulations [1, 2, 3, 4, 5] has enabled in situ monitoring of real-time dynamics and thermalization in isolated many-body quantum systems, allowing to directly probe foundational questions of quantum statistical mechanics. A powerful framework for describing thermalization in closed quantum systems is the eigenstate thermalization hypothesis (ETH) [6, 7, 8, 9, 10]. The ETH posits that in “generic” non-integrable systems – those comprising many interacting degrees of freedom – all highly-excited energy eigenstates appear “thermal” in the sense of local observable expectation values. This elucidates how local observables equilibrate thermally as the entire system evolves unitarily.
@@
However, many non-integrable systems once thought to fully obey the ETH are now understood to weakly violate it by hosting atypical eigenstates known as quantum many-body scars (QMBSs) [11, 12, 13]. Some examples of systems that weakly violate the ETH include Rydberg atom arrays [14, 15], the Heisenberg-type spin models [16, 17, 18], ultracold atoms [19, 20, 21, 22, 23, 24, 25], superconducting circuits [26, 27, 28], and numerous other [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. The ability to defy thermalization is of interest in quantum information processing [41, 42, 43], e.g., QMBSs have been used to generate Greenberger-Horne-Zeilinger (GHZ) states [44] and disorder-tunable entanglement far from equilibrium [45].
@@
A common trait of QMBS systems are long-lived, nonthermal dynamics when the system is initialized in special states with high overlap on QMBS eigenstates. This leads to phenomena such as persistent oscillations in quench dynamics and slow entanglement growth [46], both of which are unexpected in fully chaotic systems. In some models, these features have an elegant theoretical description in terms of a so-called restricted su(2) spectrum generating algebra (RSGA) [47, 48, 49, 50]. For example, the RSGA approximately describes the QMBSs in the PXP model [51, 52, 15, 53, 54, 55, 56, 57, 58, 59, 60], an effective model of Rydberg atom arrays in which persistent revivals have been observed in several experiments [14, 61, 23, 62]. On the other hand, the theoretical landscape of QMBS models is much richer due to a multitude of exact constructions of non-thermalizing states and dynamics within chaotic many-body Hamiltonians [63, 64, 65, 66, 67, 68, 69].
@@
Despite much progress in the understanding of weak ETH violations, one key practical challenge remains: for a general model described by some unfamiliar Hamiltonian, how does one verify the existence of QMBSs or other kinds of non-thermalizing states? Such a task is non-trivial as QMBSs are typically associated with chaotic systems that lack integrability or conserved quantities, making them hard to treat analytically. In models with special structure, such as the PXP model, the QMBS eigenstates stand out as outliers with the lowest entanglement entropy in the spectrum [70] and the initial states associated with them can be inferred based on semiclassical intuition [53, 71, 72]. However, this does not easily generalize to other models, including the PXP-type models with longer-range constraints [73]. Similarly, numerical tools such as exact diagonalization or tensor network methods, are limited by system size or entanglement, and often rely on model-specific insights, which restricts their applicability in general.
@@
In this work, we introduce the ScarFinder method which uncovers scar-like dynamics in general chaotic systems, without requiring prior knowledge about the model or its scar states. The key principle behind ScarFinder is that scarred dynamics are typically confined to a low-entanglement submanifold of the Hilbert space. This allows to systematically search for scar trajectories by iteratively evolving and projecting states within a suitable variational manifold, such as matrix product states (MPS) with fixed bond dimension. The only required inputs for ScarFinder are the system’s Hamiltonian and a choice of this variational manifold that is expected to support low-entanglement dynamics. Crucially, this approach does not rely on any knowledge of the scar eigenstates or their algebraic properties. Instead, the algorithm exploits the dynamical separation between scarred and thermal components during time evolution: while thermal components spread rapidly and exit the manifold, the scar components remain coherent and confined to the manifold, see Fig. 1. By combining short-time evolution with projection back to the manifold, ScarFinder effectively suppresses thermal contributions and converges toward periodic trajectories.
@@
The remainder of this paper is organized as follows. In Sec. II we motivate the ScarFinder method and outline the structure of the algorithm. In Sec. III, to facilitate understanding, we validate the algorithm using an analytically-tractable example of the spin-1 XY model that hosts exact QMBS states [17]. We demonstrate that our method recovers the known scar trajectories without any input of their algebraic construction, and we discuss in detail the convergence of the method. In Sec. IV, we apply the method to the PXP model, where we identify a new scar trajectory that enhances the previously known trajectory associated with the Néel initial state and results in robust – almost perfect – revivals in the thermodynamic limit. This highlights the power of the method both in discovering scars as well as optimizing known scar dynamics in systems without exact solutions. Finally, in Sec. V we show that ScarFinder can reveal regions of regular dynamics within “mixed phase space” of many-body quantum systems by projecting their evolution onto a low-dimensional variational manifold, where classical structures like Kolmogorov-Arnold-Moser (KAM) tori and Poincaré sections can be directly identified with high efficiency. Our conclusions are presented in Sec. VI, while Appendices contain further technical details of the algorithm, its pedagogical demonstration, and an application beyond QMBS systems, where we successfully reproduce and generalize the findings of Refs. [74, 75] for the mixed field Ising model.
@@ II.1 The algorithm and its motivation
The core idea of ScarFinder is illustrated schematically in Fig. 1. Suppose that for a given many-body Hamiltonian $\hat{H}$, there exists a variational manifold $\mathcal{M}$ that captures the scar dynamics (either exactly or to a good approximation). Specifically, let us assume there exists a state $|\psi_{0}\rangle\in\mathcal{M}$ whose time evolution remains confined to the manifold and traces out a periodic trajectory $|\psi_{0}(\theta)\rangle$, parametrized by a phase angle $\theta$.
Consider now a state slightly perturbed away from the ideal scar trajectory. As illustrated in Fig. 1(a), the state can be written as a superposition of a generic (non-scar) component $|\phi\rangle$ and a scar component $|\psi_{0}\rangle$:
@@
Figure 1: Schematic illustration of a single step of the ScarFinder algorithm. (a) The initial state $|\psi(0)\rangle$, lying on a low-entanglement variational manifold $\mathcal{M}$, is assumed to be a superposition of a scar component $|\psi_{0}\rangle$, which evolves periodically (dashed line), and a generic component $|\phi\rangle$. (b) At $t=0$, both components exhibit low entanglement, characterized by a small number of dominant Schmidt coefficients $\lambda_{k}$. (c) After a unitary time evolution step $\hat{U}(\Delta t)$, the state becomes $|\psi(\Delta t)\rangle$, with the superposition preserved. However, while the scar component $|\psi_{0}^{\prime}\rangle$ remains low-entangled, the generic component $|\phi^{\prime}\rangle$ becomes highly entangled and delocalized. (d) The delocalization of $|\phi^{\prime}\rangle$ leads to a broader distribution of Schmidt coefficients. This entanglement difference enables the separation of the two components in $|\psi(\Delta t)\rangle$: the leading Schmidt coefficients are dominated by the scar part, while the subleading terms largely originate from the thermal component. The projection step in ScarFinder leverages this structure to suppress thermal contributions and reinforce the scar trajectory.
After evolving the state under the Hamiltonian for a sufficiently long time $\Delta t$, we obtain:
$|\psi^{\prime}\rangle=e^{-i\hat{H}\Delta t}|\psi\rangle=\eta\,|\phi^{\prime}\rangle+\sqrt{1-\eta^{2}}|\psi_{0}(\theta^{\prime})\rangle,$ (2)
where $|\phi^{\prime}\rangle=e^{-i\hat{H}\Delta t}|\phi\rangle$. In systems hosting scars, it is natural to expect that the scar component $|\psi_{0}^{\prime}\rangle$ remains coherent and low-entangled, while the generic component $|\phi^{\prime}\rangle$ typically thermalizes and spreads across the Hilbert space, see Fig. 1(b).
To quantitatively reveal the difference between these two components, we perform the Schmidt decomposition across a bipartition $A{\cup}B$ in the middle of the system:
@@
where $\{\lambda_{k}\}$ are the Schmidt coefficients. The scar component $|\psi_{0}(\theta^{\prime})\rangle$ has a sharply concentrated set of Schmidt values with only a few dominant ones, while the thermalizing component $|\phi^{\prime}\rangle$ yields a broad distribution of smaller values, as illustrated in Fig. 1(c)-(d). As a result, even in the combined state $|\psi(t)\rangle$, the Schmidt spectrum exhibits a two-component structure: the dominant coefficients are governed by the scar component, while the tail contains contributions from the thermal component.
The natural separation of the Schmidt coefficients motivates a truncation procedure that effectively suppresses the non-scar part. This is achieved via projection $\hat{P}_{\mathcal{M}}$ back onto the manifold $\mathcal{M}$, yielding:
$\hat{P}_{\mathcal{M}}|\psi^{\prime}\rangle=\eta^{\prime}|\phi^{\prime}\rangle+\sqrt{1-(\eta^{\prime})^{2}}|\psi_{0}(\theta^{\prime})\rangle,$ (4)
where $\eta^{\prime}\leq\eta$ indicates the expected reduced weight of the thermal component. While we do not have a proof of Eq. (4) – and indeed do not expect it to be generally satisfied for an arbitrary $\hat{H}$ – in Sec. III and Appendix B we will demonstrate its validity in physical QMBS models.
In summary, we conjecture that each iteration of the evolution-projection cycle edges the state closer to the scar trajectory. This motivates an iterative structure of the full ScarFinder algorithm in (1). In the following Sec. II.2, we discuss the different steps and their implementation in detail.
@@ II.2 Implementation
Here, we highlight some important practical aspects of ScarFinder. First, the manifold $\mathcal{M}$ can be arbitrary. Throughout this work, we will focus on spin chains and cases where $\mathcal{M}$ is spanned by infinite matrix product states (iMPS) [76] with an $n$-site unit cell:
@@
where $A_{i}$ are matrices of maximum bond dimension $\chi$ and $|\bm{s}\rangle$ denotes basis states. This manifold faithfully captures low-entanglement states in the thermodynamic limit and it does not require prior knowledge of the scar structure, making it a natural, general-purpose choice.
@@
Input: An appropriate variational manifold $\mathcal{M}$, Hamiltonian $\hat{H}$, ScarFinder projection time step $\Delta t$, ScarFinder iteration steps $N_{\text{step}}$, target energy density $E_{\text{target}}$;
@@
3:      Evolve the $|\psi\rangle$ to $|\psi^{\prime}\rangle=\exp(-i\hat{H}\Delta t)|\psi\rangle$;
4:      Project $|\psi^{\prime}\rangle$ back onto $\mathcal{M}$, $|\psi^{\prime}\rangle\to\hat{P}_{\mathcal{M}}|\psi^{\prime}\rangle$;
5:     Enforce energy conservation if the energy of $P_{\mathcal{M}}|\psi^{\prime}\rangle$ deviates from $E_{\text{target}}$, and any crucial symmetries or constraints, $\hat{P}_{\mathcal{M}}|\psi^{\prime}\rangle\to|\psi^{\prime\prime}\rangle$;
6:     Update $|\psi\rangle\to|\psi^{\prime\prime}\rangle$;
@@
The time-evolution step (line 3) of the algorithm can be conveniently performed using standard methods, e.g., both infinite time-evolving block decimation (iTEBD) [77] and time-dependent variational principle (TDVP) [78, 79] can be used for MPS manifolds. These methods favor smaller $\Delta t$ due to the entanglement buildup, which increases the complexity of the simulation. On the other hand, ScarFinder naturally prefers large $\Delta t$, which allows the thermal component $|\phi^{\prime}\rangle$ to sufficiently delocalize over $\mathcal{M}$. Note that we will distinguish between $\Delta t$ and $t$: $\Delta t$ refers to the projection interval used during the ScarFinder optimization, while $t$ denotes the real-time evolution starting from the optimized initial state. Without some knowledge about the Hamiltonian, e.g., its information-scrambling capacity, it is difficult to determine optimal $\Delta t$. However, in several examples considered below, we find that the entanglement-based separation persists qualitatively for moderate values of $\Delta t$ that are commonly used in MPS-based methods for time evolution.
@@
Finally, after converging the results for one value of $\chi$, one might want to test their sensitivity to increasing $\chi$. The convergence of the results with $\chi$ is not guaranteed to be monotonic due to optimization instabilities or getting trapped in local minima. Nevertheless, in all scar models we studied in this paper, which do not admit an exact scar state within a fixed-bond-dimension MPS manifold, we found that increasing the bond dimension $\chi$ generally improved the performance of the algorithm. In the following Sec. III, we will apply the method to the analytically-tractable example of the spin-1 XY model [17], which will allow to concretely illustrate the above steps and build intuition about the inner workings of the algorithm.

### v3

@@ Abstract
Mechanisms that give rise to coherent quantum dynamics, such as quantum many-body scars, have recently attracted much interest as a way of controlling quantum chaos. However, identifying the presence of quantum scars in general many-body Hamiltonians remains an outstanding challenge. Here we introduce ScarFinder, a variational framework that reveals possible scar-like dynamics without prior knowledge of scar states or their algebraic structure, assuming only that such dynamics remain low in entanglement. By iteratively evolving and projecting states within a variational manifold, ScarFinder isolates scarred trajectories by suppressing thermal contributions. We validate the method on the analytically tractable spin-1 XY model, recovering the known scar dynamics, as well as the mixed field Ising model, where we capture and generalize the initial conditions previously associated with “weak thermalization”. We then apply the method to the PXP model of Rydberg atom arrays, finding a previously unknown trajectory with nearly-perfect revival dynamics in the thermodynamic limit. We also demonstrate that ScarFinder can efficiently identify the centers of stable islands in Poincaré sections of the mixed phase space that results from the projection of many-body quantum dynamics to a variational manifold. Our results establish ScarFinder as a powerful, model-agnostic tool for identifying and optimizing coherent dynamics in quantum many-body systems.
@@ I Introduction
Recent progress in quantum simulations [Bloch2012, Georgescu2014, Kjaergaard2020, Browaeys2020, MonroeRMP] has enabled in situ monitoring of real-time dynamics and thermalization in isolated many-body quantum systems, allowing to directly probe foundational questions of quantum statistical mechanics. A powerful framework for describing thermalization in closed quantum systems is the eigenstate thermalization hypothesis (ETH) [DeutschETH, SrednickiETH, RigolNature, dAlessio2016, Ueda2020]. The ETH posits that in “generic” non-integrable systems – those comprising many interacting degrees of freedom – all eigenstates apart from spectral edges appear “thermal” in the sense of local observable expectation values. This elucidates how local observables equilibrate thermally as the entire system evolves unitarily.
@@
However, many non-integrable systems once thought to fully obey the ETH are now understood to weakly violate it by hosting atypical eigenstates known as quantum many-body scars (QMBSs) [serbyn2021quantum, Moudgalya_2022, ChadranReview]. Some examples of systems that weakly violate the ETH include Rydberg atom arrays [bernien2017probing, turner2018weak], the Heisenberg-type spin models [Moudgalya2018_2, Schecter2019, yang2024phantom], ultracold atoms [Hudomal2020bosons, Zhao2020Optical, Desaules2021Tilted, scherg2021observing, Su2023, Adler2024, Evrard2024], superconducting circuits [zhang2023many, Guo2023Origin, larsen2024experimentalprotocolobservingsingle], and numerous other [ShiraishiMori, NeupertScars, Shibata202Onsager, McClarty2020, Lee2020Colored, Mohapatra2023, kolb2023stability, Srivatsa2023, Desaules2023Schwinger, Halimeh2023robustquantummany, Gotta2023, Dooley2024Dual, pizzi2024quantumscarsmanybodysystems]. The ability to defy thermalization is of interest in quantum information processing [Dooley2021, Desaules2022QFI, Dooley2023], e.g., QMBSs have been used to generate Greenberger-Horne-Zeilinger (GHZ) states [omran2019GHZ] and disorder-tunable entanglement far from equilibrium [Dong2023].
@@
A common trait of QMBS systems are long-lived, nonthermal dynamics when the system is initialized in special states with high overlap on QMBS eigenstates. In this work, we use the term “scar eigenstate” to refer to nonthermal eigenstates embedded in an otherwise thermal spectrum, and “scar initial state” to refer to an initial state that exhibits slow thermalization due to its large overlap with such eigenstates. Manifestations of slow thermalization include persistent oscillations in quench dynamics and slow entanglement growth [papić2021weakergodicitybreakinglens], both of which are unexpected in fully chaotic systems. In some models, these features have an elegant theoretical description in terms of a so-called restricted su(2) spectrum generating algebra (RSGA) [moudgalya2020eta, Mark2020, ODea2020Tunnels, Bull2020]. For example, the RSGA approximately describes the QMBSs in the PXP model [FendleySachdev, Lesanovsky2012, turner2018weak, Ho2019, Choi2019, Surace2020, Khemani2019, lin2019exact, mondragon2021fate, Omiya2022, ivanov2025exactarealawscareigenstates], an effective model of Rydberg atom arrays in which persistent revivals have been observed in several experiments [bernien2017probing, bluvstein2021controlling, Su2023, zhao2024observationquantumthermalizationrestricted]. On the other hand, the theoretical landscape of QMBS models is much richer due to a multitude of exact constructions of non-thermalizing states and dynamics within chaotic many-body Hamiltonians [ShiraishiMori, pakrouski2020many, Mark2020eta, ren2021quasisymmetry, ren2022deformed, Buca2023, Moudgalya2024].
@@
Despite much progress in the understanding of weak ETH violations, one key practical challenge remains: for a general model described by some unfamiliar Hamiltonian, how does one verify the existence of QMBSs or other kinds of non-thermalizing states? Such a task is non-trivial as QMBSs are typically associated with chaotic systems that lack integrability or conserved quantities, making them hard to treat analytically. In models with special structure, such as the PXP model, the QMBS eigenstates stand out as outliers with the lowest entanglement entropy in the spectrum [Turner2018PRB] and the initial states associated with them can be inferred based on semiclassical intuition [Ho2019, Michailidis2020, Daniel2023]. However, this does not easily generalize to other models, including the PXP-type models with longer-range constraints [kerschbaumer2024quantummanybodyscarspxp]. Similarly, numerical tools such as exact diagonalization or tensor network methods, are limited by system size or entanglement, which restricts their applicability in general.
@@
In this work, we introduce the ScarFinder method which uncovers scar-like dynamics in general chaotic systems. The key principle behind ScarFinder is that scarred dynamics are typically confined to a low-entanglement submanifold of the Hilbert space. This allows to systematically search for scar trajectories by iteratively evolving and projecting states within a suitable variational manifold, such as matrix product states (MPS) with fixed bond dimension. The only required inputs for ScarFinder are the system’s Hamiltonian and a choice of this variational manifold that harbors the low-entanglement dynamics. Crucially, this approach requires no prior knowledge of scar eigenstates or their algebraic structure, and depends only on the existence of a suitable low entanglement variational manifold. The algorithm then leverages the dynamical separation between scarred and thermal components that emerges during time evolution within this manifold: while thermal components spread rapidly and exit the manifold, the scar components remain coherent and confined to the manifold, see Fig. 1. By combining short-time evolution with projection back to the manifold, ScarFinder effectively suppresses thermal contributions and converges toward periodic trajectories.
@@
The remainder of this paper is organized as follows. In Sec. II we motivate the ScarFinder method and outline the structure of the algorithm. In Sec. III, to facilitate understanding, we validate the algorithm using an analytically-tractable example of the spin-1 XY model that hosts exact QMBS states [Schecter2019]. We demonstrate that our method recovers the known scar trajectories without any input of their algebraic construction, and we discuss in detail the convergence of the method. In Sec. IV, we apply the method to the PXP model, where we identify a new scar trajectory that enhances the previously known trajectory associated with the Néel initial state and results in robust – almost perfect – revivals in the thermodynamic limit. This highlights the power of the method both in discovering scars as well as optimizing known scar dynamics in systems without exact solutions. Finally, in Sec. V we show that ScarFinder can reveal regions of regular dynamics within “mixed phase space” of many-body quantum systems by projecting their evolution onto a low-dimensional variational manifold, where classical structures like Kolmogorov-Arnold-Moser (KAM) tori and Poincaré sections can be directly identified with high efficiency. Our conclusions are presented in Sec. VI, while Appendices contain further technical details of the algorithm, its pedagogical demonstration, and an application beyond QMBS systems, where we successfully reproduce and generalize the findings of Refs. [Banuls2011, LinMotrunich2017] for the mixed field Ising model.
@@ II The ScarFinder algorithm
Identifying initial states that lead to periodic or otherwise non-thermal dynamics within chaotic many-body systems is an important goal in understanding the applicability of the ETH and the mechanisms of its breakdown. Beyond QMBSs, recent examples of time crystals [PhysRevLett.109.160401, khemani2019briefhistorytimecrystals, RevModPhys.95.031001] and the quantum Mpemba effect [ares2025quantummpembaeffects, bhore2025quantummpembaeffectglobal] further highlight the importance of this task. In the former case, states exhibit periodic motion by spontaneously breaking time-translation symmetry, while in the latter the system’s relaxation dynamics depend counterintuitively on the effective temperature of the initial state. Despite their distinct origins, these phenomena share a common difficulty: how to pinpoint the special initial conditions in an exponentially large state space. Motivated by this need for a systematic method, below we introduce the ScarFinder algorithm and present an intuitive illustration of its underlying principle.
Figure 1: Schematic illustration of a single step of the ScarFinder algorithm. (a) The initial state $|\psi(0)\rangle$, lying on a low-entanglement variational manifold $\mathcal{M}$, is assumed to be a superposition of a scar component $|\psi_{0}\rangle$, which evolves periodically (dashed line), and a generic component $|\phi\rangle$. (b) After a unitary time evolution step $\hat{U}(\Delta t)$, the state becomes $|\psi(\Delta t)\rangle$, with the superposition preserved. However, while the scar component $|\psi_{0}^{\prime}\rangle$ remains low-entangled, the generic component $|\phi^{\prime}\rangle$ becomes highly entangled and delocalized. (c) At $t=0$, both components exhibit low entanglement, characterized by a small number of dominant Schmidt coefficients $\lambda_{k}$. (d) The delocalization of $|\phi^{\prime}\rangle$ leads to a broader distribution of Schmidt coefficients. This entanglement difference enables the separation of the two components in $|\psi(\Delta t)\rangle$: the leading Schmidt coefficients are dominated by the scar part, while the subleading terms largely originate from the thermal component. The projection step in ScarFinder leverages this structure to suppress thermal contributions and reinforce the scar trajectory.
@@ II.1 The algorithm and its motivation
The core idea of ScarFinder is illustrated schematically in Fig. 1. Suppose that for a given many-body Hamiltonian $\hat{H}$, there exists a variational manifold $\mathcal{M}$ that captures the scar dynamics (either exactly or to a good approximation). Specifically, let us assume there exists a state $|\psi_{0}\rangle\in\mathcal{M}$ whose time evolution remains confined to the manifold and traces out a periodic trajectory $|\psi_{0}(\theta)\rangle$, parametrized by a variable $\theta$.
Consider now a state slightly perturbed away from the ideal scar trajectory. As illustrated in Fig. 1(a), the state can be written as a superposition of a generic (non-scar) component $|\phi\rangle$ and a scar component $|\psi_{0}(\theta)\rangle$:
@@
After evolving the state under the Hamiltonian for a sufficiently long time $\Delta t$ under $\hat{U}(\Delta t)=\exp(-i\hat{H}\Delta t)$, we obtain:
$|\psi(\Delta t)\rangle=\eta\,\hat{U}(\Delta t)|\phi\rangle+\sqrt{1-\eta^{2}}\,\hat{U}(\Delta t)|\psi_{0}\rangle,$ (2)
In scarred systems, it is natural to expect that the second, component in the above equation remains coherent and low-entangled, while the generic component, $\hat{U}(\Delta t)|\phi\rangle$, typically thermalizes and spreads across the Hilbert space. Thus, after projection $\hat{P}_{\mathcal{M}}$ back onto the manifold $\mathcal{M}$, we obtain:
$\hat{P}_{\mathcal{M}}|\psi(\Delta t)\rangle=\eta^{\prime}\,|\phi^{\prime}\rangle+\sqrt{1-(\eta^{\prime})^{2}}\,|\psi_{0}(\theta^{\prime})\rangle,$ (3)
where the prime denotes components projected to $\mathcal{M}$ and $\eta^{\prime}\leq\eta$ indicates the expected reduced weight of the thermal component, see Fig. 1(b). While we do not have a proof of Eq. (3) – and indeed do not expect it to be generally satisfied for an arbitrary $\hat{H}$ – in Sec. III and Appendix B we will demonstrate its validity in physical QMBS models.
The difference between the two components in Eq. (3) can be revealed by performing the Schmidt decomposition across a bipartition $A{\cup}B$ in the middle of the system:
@@
where $\{\lambda_{k}\}$ are the Schmidt coefficients. Applying Eq. (4) to the scar component $|\psi_{0}(\theta^{\prime})\rangle$ gives a sharply concentrated set of Schmidt values with only a few dominant ones, while the thermalizing component $|\phi^{\prime}\rangle$ yields a broad distribution of smaller Schmidt values, as illustrated in Fig. 1(c)-(d). As a result, even in the combined state $|\psi(\Delta t)\rangle$, the Schmidt spectrum exhibits a ‘two-component’ structure: the dominant coefficients are governed by the scar component, while the tail contains contributions from the thermal component.
In summary, we conjecture that each iteration of the evolution-projection cycle, implemented via repeated applications of Eqs. (2)-(3), edges the state closer to the scar trajectory. The natural separation of the Schmidt coefficients motivates a truncation procedure that effectively suppresses the non-scar part, with the structure of the full ScarFinder presented in Algorithm 1. In the following Sec. II.2, we discuss the different steps and their implementation in detail.
@@ II.2 Implementation
Here, we highlight some important practical aspects of ScarFinder. First, the manifold $\mathcal{M}$ can be arbitrary. Throughout this work, we will focus on spin chains and cases where $\mathcal{M}$ is spanned by infinite matrix product states (iMPS) [CiracRMP] with an $n$-site unit cell:
@@
where $A_{i}$ are matrices of maximum bond dimension $\chi$ and $|\bm{s}\rangle$ denotes basis states. This manifold faithfully captures low-entanglement states in the thermodynamic limit. Provided that the target dynamics is captured by such states, we shall not require any further knowledge of the scar structure.
@@
Input: An appropriate variational manifold $\mathcal{M}$, Hamiltonian $\hat{H}$, projection time step $\Delta t$, number of iteration steps $N_{\text{step}}$, target energy density $E_{\text{target}}$;
@@
3:   Evolve the $|\psi\rangle$ to $|\psi(\Delta t)\rangle\coloneq\exp(-i\hat{H}\Delta t)|\psi\rangle$;
4:   Project back onto $\mathcal{M}$, $|\psi^{\prime}\rangle\coloneq\hat{P}_{\mathcal{M}}|\psi(\Delta t)\rangle$;
5:  Enforce energy conservation if the energy of $|\psi^{\prime}\rangle$ deviates from $E_{\text{target}}$, and any crucial symmetries or constraints, resulting in $|\psi^{\prime\prime}\rangle$;
6:  Update $|\psi\rangle\coloneq|\psi^{\prime\prime}\rangle$;
@@
The time-evolution step (line 3) of the algorithm can be conveniently performed using standard methods, e.g., both infinite time-evolving block decimation (iTEBD) [VidalTEBD] and time-dependent variational principle (TDVP) [haegeman2011time, Haegeman2014geometry] can be used for MPS manifolds. These methods favor smaller $\Delta t$ due to the entanglement buildup, which increases the complexity of the simulation. On the other hand, ScarFinder naturally prefers large $\Delta t$, which allows the thermal component $|\phi^{\prime}\rangle$ to sufficiently delocalize over $\mathcal{M}$. Note that we will distinguish between $\Delta t$ and $t$: $\Delta t$ refers to the projection interval used during the ScarFinder optimization, while $t$ denotes the real-time evolution starting from the optimized initial state. In general, larger values of $\Delta t$ enhance the separation between thermal and scar components during time evolution. Thus, one would choose $\Delta t$ to be as large as possible, given the available computational resources. In several examples considered below, we find that the entanglement-based separation persists qualitatively for moderate values of $\Delta t$ that are commonly used in MPS-based methods for time evolution.
@@
Finally, after converging the results for one value of $\chi$, one might want to test their sensitivity to increasing $\chi$. The convergence of the results with $\chi$ is not guaranteed to be monotonic due to optimization instabilities or getting trapped in local minima. Nevertheless, in all scar models we studied in this paper, which do not admit an exact scar state within a fixed-bond-dimension MPS manifold, we found that increasing the bond dimension $\chi$ generally improved the performance of the algorithm. In the following Sec. III, we will apply the method to the analytically-tractable example of the spin-1 XY model [Schecter2019], which will allow to concretely illustrate the above steps and build intuition about the inner workings of the algorithm.